# Demystifying amortized causal discovery with transformers

### v2

@@ Appendix D Further experiments
In this section, we provide additional experiments on real-world data, on the scaling properties of CSIvA in the number of training samples, and benchmark CSIvA performance in comparison to several well-established or state-of-the-art methods for causal discovery, with identifiability guarantees under different assumptions: DirectLiNGAM (Shimizu et al., 2011) for inference on linear non-Gaussian models, CAM (Bühlmann et al., 2014) for inference of additive noise models with additive mechanisms, NoGAM (Montagna et al., 2023c) and GraNDAG (Lachapelle et al., 2020) for inference on ANMs (the latter is taken from the continuous optimization literature of causal discovery, already mentioned in the related works Section 2).
@@ D.1 Experiments on real-world datasets
We consider the accuracy of CSIvA trained on different dataset configurations and tested on real-world datasets. In particular, we perform evaluation on the Tübingen pairs dataset (Mooij et al., 2016), the Sachs biological dataset (Sachs et al., 2005), the AutoMPG dataset on cars fuel consumption (Bache & Lichman, 2013) and the Sprinkler dataset, a simple dataset on the causal relations between the binary categorical variables rain, sprinkler on/off, wet grass. Given that our algorithms are trained on bivariate models, from each multivariate dataset we extract all possible two variables subgraphs where this operation does not introduce new confounding effects. This results in $9$ datasets from Sachs, $3$ datasets from AutoMPG, and $2$ datasets from Sprinkler. We consider $102$ pairs from the Tübingen dataset.
@@
In Fig. 6 we illustrate the average accuracy per dataset type (Sachs, AutoMPG, Sprinkler, Tübingen) of each CSIvA model. In particular, we want to probe the goodness of mixed training in real-world scenarios. To this end, we train four architectures on the following dataset configurations: linear-mlp, anm-mlp, pnl-mlp, mixed-mixed, where the latter denotes the model trained on SCMs with linear, additive nonlinear and post-nonlinear mechanisms, and Beta, Gamma, Gumbel, Exponential, MLP, and Uniform noise distributions. We find the following interesting outcome: the mixed-mixed architecture is on par with the others on the Sprinkler and the Sachs datasets and outperforms the other methods on the AutoMPG and the Tübingen pairs datasets. Despite these results must be taken cautiously, they provide evidence of a strong result, that mixed training appears to be beneficial even in real-world scenarios, those of actual interest in applications.
@@
We probe CSIvA test generalization in comparison with DirectLiNGAM, CAM, NoGAM, and GraNDAG methods. According to Fig. 7, interestingly we find that the mixed-mixed CSIvA model (trained on SCMs with linear, additive nonlinear and post-nonlinear mechanisms, and Beta, Gamma, Gumbel, Exponential, MLP, Uniform noise distributions) matches with or outperforms the other methods on all the test tasks. This provides additional empirical evidence on the benefits of the mixed training procedure we propose to achieve better test generalization.
@@ D.2 Benchmarking CSIvA generalization with classical causal discovery algorithms
In this section, we analyze the results of Fig. 8, where we compare the CSIvA trained on mixed mechanisms (linear, nonlinear, post-nonlinear) and mixed noises (all noise except for Gaussian) with the benchmark methods DirectLiNGAM, CAM, NoGAM, GraNDAG. Given that we want to probe CSIvA test generalization, we run inference over the following dataset configurations: linear-mixed (i.e. SCMs considering all possible noise distributions, except for Gaussian), anm-mixed, pnl-mixed, and mixed-mlp (i.e. SCMs with linear, nonlinear, post-nonlinear mechanisms). Fig. 8 shows results in line with our expectations: DirectLiNGAM, CAM, NoGAM, and GraNDAG achieve their best accuracy on data generated by SCMs respecting their assumptions, while degrading their performance on the other models; the CSIvA architecture trained on a mixture of SCMs with different mechanisms and noise distributions matches with or tops all other methods, in all the considered settings (while being outperformed on the anm and linear data, CSIvA still retains good average SHD accuracy).
@@ D.3 Experiments with different sizes of the training dataset
In this section, we explore how CSIvA test generalization scales when training occurs on different numbers of training samples. In the experiments on the main manuscript, each algorithm is optimized on $15000$ datasets, where each dataset and the underlying causal graph corresponds to a training data point. We now compare the test SHD when training occurs on $5000$ and $10000$ datasets. One clear point emerges from the results of Fig. 9, that is our results on the benefits of the mixed training procedure are consistent for each size of the training dataset we considered. Moreover, we note that the performance of CSIvA does not appear to degrade due to the decrease in the number of training points.
@@ D.4 Can we learn to infer causal order from linear Gaussian data?
We ask whether CSIvA trained on non-identifiable models can implicitly learn to predict the causal direction of identifiable SCMs. For this purpose, we consider CSIvA optimized on linear Gaussian data and test its performance on several datasets sampled from structural causal models with different configurations of mechanisms and noise distributions: linear-mixed (with noise terms sampled according to all distributions except for Gaussian), anm-mixed, pnl-mixed, mixed-mlp (with mechanisms generated according to linear, nonlinear, post-nonlinear equations), anm-gauss, and pnl-gauss. The results of Fig. 10 present strong evidence that models trained on non-identifiable SCMs can not infer the causal order: in fact, we see that consistently across all datasets CSIvA average SHD approximates $0.5$, the performance of classification with a coin flip. This is in line with our expectations. In agreement with our Hypothesis 1, in Section 4.3 we have empirically shown that CSIvA can model the class of the SCM generating the observed data and exploit this information to infer the correct causal DAG (instead of a less specific Markov equivalence class) when this is identifiable. Moreover, in our Section 4.2 our experiments show that CSIvA can not generalize to SCM classes unseen during training. In light of these findings, it is intuitive that an architecture trained on non-identifiable linear Gaussian data can only try to fit a linear Gaussian model, irrespective of the input data. Then, when inferring the causal direction, given that CSIvA assumes the data to be generated according to a linear Gaussian SCM, both the forward and backward directions are equally plausible, which explains the observed SHD close to $0.5$.
@@ D.5 Experiments with bivariate independent graphs
In the main manuscript, we consider training and testing of CSIvA on bivariate graphs with an edge: $X\rightarrow Y$, $Y\rightarrow X$. This can be phrased as a classification problem with two labels. We motivate our choice by noticing that, in the bivariate setting, identifiability is a property of connected graphs: the empty graph with no edge defines a Markov equivalence class with one element, i.e. a singleton. This is known to be identifiable without explicit assumptions on the functional form of the mechanisms or the noise term distributions in the causal model. The goal of this section is to show that, if we include datasets generated according to an empty graph in the training procedure, CSIvA can learn to disambiguate between the three classes (the empty graph, $X\rightarrow Y$, $Y\rightarrow X$). To motivate our claim, we notice that classifying empty and connected graphs can be done by testing independence between the input variables: previous works phrase independence testing as a classification task and show that this can be learned via deep neural networks (Bellot & van der Schaar, 2019, Sen et al., 2017). The experiments of Fig. 11 sustain our claim. We consider three CSIvA architectures, each trained on independent pairs and one between linear, nonlinear, or post-nonlinear data. Our results show that the neural network can learn to disambiguate between the three classes in all scenarios.
Figure 11: Average SHD (the lower, the better) for CSIvA trained on independent pairs and one between linear, nonlinear, or post-nonlinear data. Each algorithm is tested on the same class of structural causal models it was trained on. We note that in all three scenarios, CSIvA learns to distinguish all classes with almost optimal accuracy (i.e., SHD close to 00.
@@ D.6 Mixed training with unlimited budget
In the main text of this paper, we compare neural networks trained on a mix of structural causal models (e.g. noise distributions, or mechanism types), to models trained on a single mechanism-noise combination, where all models have the same amount of training data, $15,000$ datasets. In mixed training, we split these evenly, so a "lin, nl" model is trained on $7,500$ datasets from linear SCMs, and $7,500$ from nonlinear SCMs. Our results in this framework are promising, and show that for many combinations of SCM types, we can train one model instead of two, and achieve good progress, while making a $50\%$ savings on training costs. However, if our training budget is high/unlimited, we should also ask whether we can we achieve the same performance as a model trained on a single SCM type. Fig. 12 shows good results in this direction - the models trained with the same number of datasets per SCM type as an unmixed model had similar (or even better, for PNL data) performance as the un-mixed model trained on the same SCM type as the test data. These mixed models are also significantly more useful than having 2 or 3 separate models per SCM type, as they have good across-the-board performance. However, if we used the same computational resources to train 3 separate networks (one for each mechanism type) and wanted to use them for causal discovery on a dataset with unknown assumptions, we would be left with the rather difficult task of deciding which model to trust.
@@
Figure 12: Mixtures of causal mechanisms, with varying amounts of training data. We train eight models on samples from structural casual models with different mechanisms. Four (in purple), were trained on $15,000$ samples for each SCM type (so the "lin,nl" model saw $30,000$ samples in total, and the "all" model saw $45,000$), and the other four (blue) are the same as in Fig. 4, and were trained on $15,000$ samples in total, evenly split between the SCM types they were trained on. We compare their test SHD (the lower, the better) against networks trained on datasets generated according to a single type of mechanism. The dashed line indicates the test SHD of a model trained on samples with the same mechanisms as the test SCM. Training on multiple causal models with different mechanisms (mixed bars) always improves performance compared to training on single SCMs.
@@ D.7 Experiments with Gaussian process nonlinear mechanisms
In this section we present results obtained training and testing CSIvA on synthetic data with nonlinear mechanisms sampled from a Gaussian process with a unit bandwidth RBF kernel (we call data generated according to this approach as GP-data). In particular, for each variable $X_{i}$ node of the graph $\mathcal{G}$ generated according to model (1) we define the nonlinear mechanism $f_{i}(X_{\operatorname{PA}_{i}^{\mathcal{G}}})=\mathcal{N}(\mathbf{0},K(X_{\operatorname{PA}_{i}^{\mathcal{G}}},X_{\operatorname{PA}_{i}^{\mathcal{G}}}))$, a multivariate normal distribution centered at zero and with covariance matrix as the Gaussian kernel $K(X_{\operatorname{PA}_{i}^{\mathcal{G}}},X_{\operatorname{PA}_{i}^{\mathcal{G}}})$, where $X_{\operatorname{PA}_{i}^{\mathcal{G}}}$ are the observations of the parents of the node $X_{i}$. Together with our strategy adopted in the experiments in the main text of parametrizing nonlinearities with random neural networks (NN-data), this is one of the most common approaches in the literature.
@@
We present an extensive list of works adopting Gaussian processes for the sampling of nonlinear mechanisms: Rolland et al. (2022), Montagna et al. (2023a; b; c), Bühlmann et al. (2014), Mooij et al. (2016), Lachapelle et al. (2020), Wang et al. (2021), Chen et al. (2023), Mooij et al. (2011), Monti et al. (2019). This suggests that our data generation strategy is established in the causality literature.
@@ Appendix E Theoretical results and proofs
In this section, we state and prove the identifiability of the post-ANM discussed in Section 4.3, as a corollary of Theorem 1 of Hoyer et al. (2008). The forward and backward models of equations (5) and (6) for the pair of random variables $X,Y$ is given by:
@@
with $f_{2},g_{2}$ invertible functions, $N_{Y},X$ independent random variables, and $N_{X},Y$ independent random variables. Equivalently, we can frame forward and backward causal models for $\tilde{X},\tilde{Y}$, as in equations (7) and (8):
@@
Let $p_{N_{Y}},h_{X},h_{Y}$ be fixed, and define $\nu_{Y}\coloneqq\log p_{N_{Y}}$, $\xi\coloneqq\log p_{\tilde{X}}$. Suppose that $p_{N_{Y}}$ and $p_{\tilde{X}}$ are strictly positive densities, and that $\nu_{Y},\xi,f_{1},f_{2},g_{1},$ and $g_{2}$ are three times differentiable. Further, assume that for a fixed pair $h_{Y},\nu_{Y}$ exists $\tilde{y}\in\mathbb{R}$ s.t. $\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h^{\prime}_{Y}(\tilde{x})\neq 0$ is satisfied for all but a countable set of points $\tilde{x}\in\mathbb{R}$. Then, the set of all densities $p_{\tilde{X}}$ of $\tilde{X}$ such that both equations (5) and (6) are satisfied is contained in a 2-dimensional space.
Before stating the proof of Proposition 1, we show under which condition the pair of random variables $X,Y$ satisfies the forward and backward models of equations (5), (6): this is relevant for our discussion, as the proof of Proposition 1 consists of showing that this condition is almost never satisfied.
@@
Theorem 2 (Theorem 1 of Zhang & Hyvärinen (2009)).
Assume that $X,Y$ satisfies both causal relations of equations (5) and (6). Further, suppose that $p_{N_{Y}}$ and $p_{\tilde{X}}$ are positive densities on the support of $N_{Y}$ and $\tilde{X}$ respectively, and that $\nu_{Y},\xi,f_{1},f_{2},g_{1},$ and $g_{2}$ are third order differentiable. Then, for each pair $(\tilde{x},\tilde{y})$ satisfying $\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h_{Y}(\tilde{x})\neq 0$, the following differential equation holds:
@@
Given that equations (5) and (6) hold, this implies that the forward and backward models on $\tilde{X},\tilde{Y}$ of equations (7) and (8) are also valid, namely that:
@@
$p_{{\tilde{X},\tilde{Y}}}(\tilde{x},\tilde{y})=p_{{\tilde{Y}|\tilde{X}}}(\tilde{y}|\tilde{x})p_{{\tilde{X}}}(\tilde{x})=p_{{N_{Y}}}(\tilde{y}-h_{Y}(\tilde{x}))p_{{\tilde{X}}}(\tilde{x}),$
@@
$\pi(\tilde{x},\tilde{y})=\nu_{X}(\tilde{x}-h_{X}(\tilde{y}))+\eta(\tilde{y}).$ (12)
From (12), we have that:
$\frac{\partial^{2}}{\partial\tilde{x}^{2}}\pi(\tilde{x},\tilde{y})=\nu^{\prime\prime}_{X}(\tilde{x}-h_{X}(\tilde{y}))$
$\frac{\partial^{2}}{\partial\tilde{x}\partial\tilde{y}}\pi(\tilde{x},\tilde{y})=-\nu^{\prime\prime}_{X}(\tilde{x}-h_{X}(\tilde{y}))h^{\prime}_{X}(\tilde{y}),$
@@
Computing the same set of partial derivatives from (11), we find:
$\frac{\partial^{2}}{\partial\tilde{x}^{2}}\pi(\tilde{x},\tilde{y})=\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))(h^{\prime}_{Y}(\tilde{x}))^{2}-\nu^{\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h^{\prime\prime}_{Y}(\tilde{x})+\xi^{\prime\prime}(\tilde{x})$
$\frac{\partial^{2}}{\partial\tilde{x}\partial\tilde{y}}\pi(\tilde{x},\tilde{y})=-\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h^{\prime}_{Y}(\tilde{x}).$
@@
where we drop the input arguments for conciseness. The equality with 00 is given by the equality with (13). Manipulating the above expression, the first claim follows.
@@
for any $(\tilde{y},n_{x})$. According to equations (7), (8), we have that:
@@
which must be equal to zero, being equal to the LHS of (14). Thus, we conclude that

### v3

@@ Appendix D Further experiments
In this section, we provide additional experiments on real-world data, on the scaling properties of CSIvA in the number of training samples, and benchmark CSIvA performance in comparison to several well-established or state-of-the-art methods for causal discovery, with identifiability guarantees under different assumptions: DirectLiNGAM (Shimizu et al., 2011) for inference on linear non-Gaussian models, CAM (Bühlmann et al., 2014) for inference of additive noise models with additive mechanisms, NoGAM (Montagna et al., 2023c) and GraNDAG (Lachapelle et al., 2020) for inference on ANMs (the latter is taken from the continuous optimization literature of causal discovery, already mentioned in the related works Section˜2).
@@ D.1 Experiments on real-world datasets
We consider the accuracy of CSIvA trained on different dataset configurations and tested on real-world datasets. In particular, we perform evaluation on the Tübingen pairs dataset (Mooij et al., 2016), the Sachs biological dataset (Sachs et al., 2005), the AutoMPG dataset on cars fuel consumption (Bache and Lichman, 2013) and the Sprinkler dataset, a simple dataset on the causal relations between the binary categorical variables rain, sprinkler on/off, wet grass. Given that our algorithms are trained on bivariate models, from each multivariate dataset we extract all possible two variables subgraphs where this operation does not introduce new confounding effects. This results in $9$ datasets from Sachs, $3$ datasets from AutoMPG, and $2$ datasets from Sprinkler. We consider $102$ pairs from the Tübingen dataset.
@@
In Fig.˜6 we illustrate the average accuracy per dataset type (Sachs, AutoMPG, Sprinkler, Tübingen) of each CSIvA model. In particular, we want to probe the goodness of mixed training in real-world scenarios. To this end, we train four architectures on the following dataset configurations: linear-mlp, anm-mlp, pnl-mlp, mixed-mixed, where the latter denotes the model trained on SCMs with linear, additive nonlinear and post-nonlinear mechanisms, and Beta, Gamma, Gumbel, Exponential, MLP, and Uniform noise distributions. We find the following interesting outcome: the mixed-mixed architecture is on par with the others on the Sprinkler and the Sachs datasets and outperforms the other methods on the AutoMPG and the Tübingen pairs datasets. Despite these results must be taken cautiously, they provide evidence of a strong result, that mixed training appears to be beneficial even in real-world scenarios, those of actual interest in applications.
@@
We probe CSIvA test generalization in comparison with DirectLiNGAM, CAM, NoGAM, and GraNDAG methods. According to Fig.˜7, interestingly we find that the mixed-mixed CSIvA model (trained on SCMs with linear, additive nonlinear and post-nonlinear mechanisms, and Beta, Gamma, Gumbel, Exponential, MLP, Uniform noise distributions) matches with or outperforms the other methods on all the test tasks. This provides additional empirical evidence on the benefits of the mixed training procedure we propose to achieve better test generalization.
@@ D.2 Benchmarking CSIvA generalization with classical causal discovery algorithms
In this section, we analyze the results of Fig.˜8, where we compare the CSIvA trained on mixed mechanisms (linear, nonlinear, post-nonlinear) and mixed noises (all noise except for Gaussian) with the benchmark methods DirectLiNGAM, CAM, NoGAM, GraNDAG. Given that we want to probe CSIvA test generalization, we run inference over the following dataset configurations: linear-mixed (i.e. SCMs considering all possible noise distributions, except for Gaussian), anm-mixed, pnl-mixed, and mixed-mlp (i.e. SCMs with linear, nonlinear, post-nonlinear mechanisms). Fig.˜8 shows results in line with our expectations: DirectLiNGAM, CAM, NoGAM, and GraNDAG achieve their best accuracy on data generated by SCMs respecting their assumptions, while degrading their performance on the other models; the CSIvA architecture trained on a mixture of SCMs with different mechanisms and noise distributions matches with or tops all other methods, in all the considered settings (while being outperformed on the anm and linear data, CSIvA still retains good average SHD accuracy).
@@ D.3 Experiments with different sizes of the training dataset
In this section, we explore how CSIvA test generalization scales when training occurs on different numbers of training samples. In the experiments on the main manuscript, each algorithm is optimized on $15000$ datasets, where each dataset and the underlying causal graph corresponds to a training data point. We now compare the test SHD when training occurs on $5000$ and $10000$ datasets. One clear point emerges from the results of Fig.˜9, that is our results on the benefits of the mixed training procedure are consistent for each size of the training dataset we considered. Moreover, we note that the performance of CSIvA does not appear to degrade due to the decrease in the number of training points.
@@ D.4 Can we learn to infer causal order from linear Gaussian data?
We ask whether CSIvA trained on non-identifiable models can implicitly learn to predict the causal direction of identifiable SCMs. For this purpose, we consider CSIvA optimized on linear Gaussian data and test its performance on several datasets sampled from structural causal models with different configurations of mechanisms and noise distributions: linear-mixed (with noise terms sampled according to all distributions except for Gaussian), anm-mixed, pnl-mixed, mixed-mlp (with mechanisms generated according to linear, nonlinear, post-nonlinear equations), anm-gauss, and pnl-gauss. The results of Fig.˜10 present strong evidence that models trained on non-identifiable SCMs can not infer the causal order: in fact, we see that consistently across all datasets CSIvA average SHD approximates $0.5$, the performance of classification with a coin flip. This is in line with our expectations. In agreement with our motivating hypothesis (Section˜4), in Section˜4.3 we have empirically shown that CSIvA can model the class of the SCM generating the observed data and exploit this information to infer the correct causal DAG (instead of a less specific Markov equivalence class) when this is identifiable. Moreover, in our Section 4.2 our experiments show that CSIvA can not generalize to SCM classes unseen during training. In light of these findings, it is intuitive that an architecture trained on non-identifiable linear Gaussian data can only try to fit a linear Gaussian model, irrespective of the input data. Then, when inferring the causal direction, given that CSIvA assumes the data to be generated according to a linear Gaussian SCM, both the forward and backward directions are equally plausible, which explains the observed SHD close to $0.5$.
@@ D.5 Experiments with bivariate independent graphs
In the main manuscript, we consider training and testing of CSIvA on bivariate graphs with an edge: $X\rightarrow Y$, $Y\rightarrow X$. This can be phrased as a classification problem with two labels. We motivate our choice by noticing that, in the bivariate setting, identifiability is a property of connected graphs: the empty graph with no edge defines a Markov equivalence class with one element, i.e. a singleton. This is known to be identifiable without explicit assumptions on the functional form of the mechanisms or the noise term distributions in the causal model. The goal of this section is to show that, if we include datasets generated according to an empty graph in the training procedure, CSIvA can learn to disambiguate between the three classes (the empty graph, $X\rightarrow Y$, $Y\rightarrow X$). To motivate our claim, we notice that classifying empty and connected graphs can be done by testing independence between the input variables: previous works phrase independence testing as a classification task and show that this can be learned via deep neural networks (Bellot and van der Schaar, 2019; Sen et al., 2017). The experiments of Fig.˜11 sustain our claim. We consider three CSIvA architectures, each trained on independent pairs and one between linear, nonlinear, or post-nonlinear data. Our results show that the neural network can learn to disambiguate between the three classes in all scenarios.
Figure 11: Average SHD (the lower, the better) for CSIvA trained on independent pairs and one between linear, nonlinear, or post-nonlinear data. Each algorithm is tested on the same class of structural causal models it was trained on. We note that in all three scenarios, CSIvA learns to distinguish all classes with almost optimal accuracy (i.e., SHD close to 0.
@@ D.6 Mixed training with unlimited budget
In the main text of this paper, we compare neural networks trained on a mix of structural causal models (e.g. noise distributions, or mechanism types), to models trained on a single mechanism-noise combination, where all models have the same amount of training data, $15,000$ datasets. In mixed training, we split these evenly, so a "lin, nl" model is trained on $7,500$ datasets from linear SCMs, and $7,500$ from nonlinear SCMs. Our results in this framework are promising, and show that for many combinations of SCM types, we can train one model instead of two, and achieve good progress, while making a $50\%$ savings on training costs. However, if our training budget is high/unlimited, we should also ask whether we can we achieve the same performance as a model trained on a single SCM type. Fig.˜12 shows good results in this direction - the models trained with the same number of datasets per SCM type as an unmixed model had similar (or even better, for PNL data) performance as the un-mixed model trained on the same SCM type as the test data. These mixed models are also significantly more useful than having 2 or 3 separate models per SCM type, as they have good across-the-board performance. However, if we used the same computational resources to train 3 separate networks (one for each mechanism type) and wanted to use them for causal discovery on a dataset with unknown assumptions, we would be left with the rather difficult task of deciding which model to trust.
@@
Figure 12: Mixtures of causal mechanisms, with varying amounts of training data. We train eight models on samples from structural casual models with different mechanisms. Four (in purple), were trained on $15,000$ samples for each SCM type (so the "lin,nl" model saw $30,000$ samples in total, and the "all" model saw $45,000$), and the other four (blue) are the same as in Fig.˜4, and were trained on $15,000$ samples in total, evenly split between the SCM types they were trained on. We compare their test SHD (the lower, the better) against networks trained on datasets generated according to a single type of mechanism. The dashed line indicates the test SHD of a model trained on samples with the same mechanisms as the test SCM. Training on multiple causal models with different mechanisms (mixed bars) always improves performance compared to training on single SCMs.
@@ D.7 Experiments with Gaussian process nonlinear mechanisms
In this section we present results obtained training and testing CSIvA on synthetic data with nonlinear mechanisms sampled from a Gaussian process with a unit bandwidth RBF kernel (we call data generated according to this approach as GP-data). In particular, for each variable $X_{i}$ node of the graph $\mathcal{G}$ generated according to model equation 1 we define the nonlinear mechanism $f_{i}(X_{\operatorname{PA}_{i}^{\mathcal{G}}})=\mathcal{N}(\mathbf{0},K(X_{\operatorname{PA}_{i}^{\mathcal{G}}},X_{\operatorname{PA}_{i}^{\mathcal{G}}}))$, a multivariate normal distribution centered at zero and with covariance matrix as the Gaussian kernel $K(X_{\operatorname{PA}_{i}^{\mathcal{G}}},X_{\operatorname{PA}_{i}^{\mathcal{G}}})$, where $X_{\operatorname{PA}_{i}^{\mathcal{G}}}$ are the observations of the parents of the node $X_{i}$. Together with our strategy adopted in the experiments in the main text of parametrizing nonlinearities with random neural networks (NN-data), this is one of the most common approaches in the literature.
@@
We present an extensive list of works adopting Gaussian processes for the sampling of nonlinear mechanisms: Rolland et al. (2022); Montagna et al. (2023a; b; c); Bühlmann et al. (2014); Mooij et al. (2016); Lachapelle et al. (2020); Wang et al. (2021); Chen et al. (2023); Mooij et al. (2011); Monti et al. (2019). This suggests that our data generation strategy is established in the causality literature.
@@ Appendix E Theoretical results and proofs
In this section, we state and prove the identifiability of the post-ANM discussed in Section˜4.3, as a corollary of Theorem 1 of Hoyer et al. (2008). The forward and backward models of equations equation 5 and equation 6 for the pair of random variables $X,Y$ is given by:
@@
with $f_{2},g_{2}$ invertible functions, $N_{Y},X$ independent random variables, and $N_{X},Y$ independent random variables. Equivalently, we can frame forward and backward causal models for $\tilde{X},\tilde{Y}$, as in equations equation 7 and equation 8:
@@
Let $p_{N_{Y}},h_{X},h_{Y}$ be fixed, and define $\nu_{Y}\coloneqq\log p_{N_{Y}}$, $\xi\coloneqq\log p_{\tilde{X}}$. Suppose that $p_{N_{Y}}$ and $p_{\tilde{X}}$ are strictly positive densities, and that $\nu_{Y},\xi,f_{1},f_{2},g_{1},$ and $g_{2}$ are three times differentiable. Further, assume that for a fixed pair $h_{Y},\nu_{Y}$ exists $\tilde{y}\in\mathbb{R}$ s.t. $\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h^{\prime}_{Y}(\tilde{x})\neq 0$ is satisfied for all but a countable set of points $\tilde{x}\in\mathbb{R}$. Then, the set of all densities $p_{\tilde{X}}$ of $\tilde{X}$ such that both equations equation 5 and equation 6 are satisfied is contained in a 2-dimensional space.
Before stating the proof of Proposition 1, we show under which condition the pair of random variables $X,Y$ satisfies the forward and backward models of equations equation 5, equation 6: this is relevant for our discussion, as the proof of Proposition 1 consists of showing that this condition is almost never satisfied.
@@
Theorem 2 (Theorem 1 of Zhang and Hyvärinen (2009)).
[…]
@@
Given that equations equation 5 and equation 6 hold, this implies that the forward and backward models on $\tilde{X},\tilde{Y}$ of equations equation 7 and equation 8 are also valid, namely that:
@@
[…]
@@
[…]
From equation 12, we have that:
$)\displaystyle\frac{\partial^{2}}{\partial\tilde{x}^{2}}\pi(\tilde{x},\tilde{y})=\nu^{\prime\prime}_{X}(\tilde{x}-h_{X}(\tilde{y}$))
[…]
@@
Computing the same set of partial derivatives from equation 11, we find:
[…]
@@
where we drop the input arguments for conciseness. The equality with 0 is given by the equality with equation 13. Manipulating the above expression, the first claim follows.
@@
for any $(\tilde{y},n_{x})$. According to equations equation 7, equation 8, we have that:
@@
which must be equal to zero, being equal to the LHS of equation 14. Thus, we conclude that